OpenStudy (anonymous):

If a family has 5 children, find the probability that at least one child is boy. Can you explain how to solve for the answer?

OpenStudy (anonymous):

n=5 let x be the probability that the child is a boy p(x>=1)=[(1/5)(1-1/5)+(2/5)(1-2/5)+(3/5)(1-3/5)+(4/5)(1-4/5)+(5/5)(1-5/5)] =0.52 that's the probability that the family might be having a boy child

OpenStudy (anonymous):

Only way there can't be a boy is if there are all girls. Chance of all girls is 1/32 so at least 1 boy is 1 - 1/32 =31/32 (I think)

OpenStudy (anonymous):

do u mean the chance is approximately 97%?

OpenStudy (anonymous):

Or 3% of them all being girls....